# LRFusionPR: A Polar BEV-Based LiDAR-Radar Fusion Network for Place Recognition

### v2

@@ IV-G Study on Viewpoint Rotation Invariance
In this section, we evaluate the viewpoint rotation invariance property of our proposed LRFusionPR. Following [6, 9], we conduct experiments by rotating each query sample by $\left[60\textdegree,120\textdegree,180\textdegree,240\textdegree,300\textdegree\right]$. Fig. 6 illustrates the rotation robustness of BEV-based baselines, LRFusionPR, and its variants on the SQ split. AutoPlace [4] and LRFusionPR-Cart, both using Cartesian BEV as input, exhibit significant viewpoint sensitivity. LRFusionPR-Cart-Aug mitigates this issue through data augmentation, but does not achieve complete rotation invariance. BEVPlace++ [7] achieves rotation invariance via the devised network architecture combined with data augmentation. In contrast, our LRFusionPR does not require specially designed modules. By exploiting the properties of polar BEV, rotation invariance can be achieved with only data augmentation.
TABLE VII: Runtime analysis and comparision
Method Extraction Latency [ms] Retrieval Latency [ms] Total [ms] GFLOPs
@@
LRFusionPR (ours) 6.53 0.25 6.78 7.22
IV-H Efficiency Analysis
In this section, we provide an efficiency analysis of our LRFusionPR. We compare the runtime and computational complexity of LRFusionPR with the baselines. We report the average time of descriptor extraction and similarity retrieval on the BS split. Specifically, the descriptor extraction measures the time to generate a descriptor with batch size 1. The similarity retrieval measures the time to retrieve the top-1 candidate from the database. The FAISS library [34] is used to accelerate retrieval. As shown in Tab. VII, LRFusionPR only takes 6.78 ms to perform a place recognition. Furthermore, generating polar BEVs takes only about 7 ms. This allows our method to operate at a frequency above 70 Hz, demonstrating its real-time capability. Additionally, our LRFusionPR achieves low memory consumption and computational complexity with 24.4 M parameters and 7.22 GFLOPs.
@@ V Conclusion
In this paper, we propose a LiDAR-radar fusion place recognition network that is applicable to both single-chip radar and scanning radar. We first harmonize LiDAR and radar data into polar BEV representations, thereby bridging the domain gap between sensors. Furthermore, we propose a devised dual-branch network. Specifically, by integrating cross-modality attention and distillation, we achieve robust and accurate descriptor extraction. Experiments on three datasets show that our LiDAR-radar fusion strategy outperforms all LiDAR-only, radar-only, and LiDAR-camera fusion baselines in terms of recognition accuracy. In addition, we demonstrate that the judicious integration of radar can achieve superior weather robustness through simulated foggy weather experiments. This robustness is difficult to achieve with LiDAR-only or LiDAR-camera fusion approaches.

### v3

@@ IV-F Study on Viewpoint Rotation Invariance
In this section, we evaluate the viewpoint rotation invariance property of our proposed LRFusionPR. Following [6, 9], we conduct experiments by rotating each query sample by $\left[60\textdegree,120\textdegree,180\textdegree,240\textdegree,300\textdegree\right]$. Fig. 6 illustrates the rotation robustness of BEV-based baselines, LRFusionPR, and its variants on the SQ split. AutoPlace [4] and LRFusionPR-Cart, both using Cartesian BEV as input, exhibit significant viewpoint sensitivity. LRFusionPR-Cart-Aug mitigates this issue through data augmentation, but does not achieve complete rotation invariance. BEVPlace++ [7] achieves rotation invariance via the computationally expensive group convolution architecture combined with data augmentation. In contrast, leveraging the properties of polar BEV, our LRFusionPR achieves strong viewpoint rotation robustness with only simple data augmentation, eliminating the need for specially designed modules. This validates our approach of exploiting polar BEV’s inherent characteristics for rotation-invariant place recognition.
TABLE VII: Ablation study on multimodal training
Session LRFusionPR-R-S LRFusionPR-R-M
AR@1 AR@5 AR@10 AR@1 AR@5 AR@10
BS 50.60 63.77 69.21 61.44 74.43 78.68
SON 56.76 70.78 75.34 61.15 76.52 83.36
SQ 69.33 82.61 86.84 78.38 91.54 94.48
Sejong 92.39 96.45 97.39 93.26 96.81 97.83
Riverside 94.08 98.03 98.55 94.78 98.67 99.30
Bor-Snow 98.15 98.78 99.12 98.54 99.19 99.45
TABLE VIII: Runtime analysis and comparison
Method Extraction [ms] Retrieval [ms] Total [ms] GFLOPs
@@
LRFusionPR 6.53 0.25 6.78 7.22
GFLOPS for sparse convolution-based baselines are not reported due to unmeasurable complexity.
IV-G Efficiency Analysis
In this section, we provide an efficiency analysis of our LRFusionPR. We compare the runtime and computational complexity of LRFusionPR with the baselines. We report the average time of descriptor extraction and similarity retrieval on the BS split. As shown in Tab. VIII, LRFusionPR only takes 6.78 ms to perform a place recognition , demonstrating its real-time capability. Additionally, our LRFusionPR achieves low memory consumption and computational complexity with 24.4 M parameters and 7.22 GFLOPs.
@@ V Conclusion
In this paper, we propose a LiDAR-radar fusion place recognition network applicable to both single-chip and scanning radar. We unify multi-sensor data into polar BEVs, eliminating domain discrepancies while enabling rotation-invariant place recognition. Furthermore, our proposed dual-branch network then leverages deep attentive fusion and cross-modality self-distillation to extract discriminative representations from noisy radar data. Validated across four datasets and diverse weather conditions, our LRFusionPR consistently outperforms all baselines, demonstrating multimodal complementarity for accurate and robust recognition.